OpenStudy (anonymous):

How do I find the slope of (-4,-6) and (-5,-1)?

OpenStudy (chriss):

\[slope=\frac{y_{1}-y_{0}}{x_{1}-x_{0}}\] \[\frac{-1-(-6)}{-5-(-4)}\] \[\frac{5}{-1}\] m=-5